OpenStudy (anonymous):

Which of the following recursive formulas represent the same arithmetic sequence as the explicit formula an=5+(n-1)2?

OpenStudy (anonymous):

where r the options?

OpenStudy (amoodarya):

to find choice find \[a_n-a_{n-1}\]

OpenStudy (anonymous):

(An-1+5)•2 An=an-1•5 An=an-1+5 An=an-1+2

OpenStudy (welshfella):

the common difference is 2 each term is obtained from the last by adding 2

OpenStudy (welshfella):

which of the options will fit that?

OpenStudy (anonymous):

a?

OpenStudy (amoodarya):

\[a_n=5+2(n-1)\\a_{n-1}=5+2((n-1)-1)=5+2(n-2)\\a_n-a_{n-1}=\left[ 5+2(n-1) \right]-\left[ 5+2(n-2) \right]=\\5-5+2(n-1-(n-2))\\a_n-a_{n-1}=2\]

OpenStudy (welshfella):

no not a

OpenStudy (anonymous):

Oh! I thinks it's b!

OpenStudy (welshfella):

i think you are guessing the first term is 5 and the second is 7 put an = 7 and an-1 = 5 and see which of the options fit

OpenStudy (anonymous):

An-1=5

OpenStudy (welshfella):

try option b An = 7 An-1 * 5 = 5*5 = 27 that does not fit

OpenStudy (welshfella):

now try c and d